# Haptic bilateral teleoperation system for free-hand dental procedures

### v1

@@ III-B2 Human tremors filtering
III-B3 Eye-hand coordination
In an HBTS, coupling two robots with different kinematics requires the use of mapping and scaling techniques. The literature shows that this coupling can be achieved by converting the motion of the HD into that of the FR through simple kinematic transformations, and appropriately scaling it to either expand the workspace size of the HD or enhance control accuracy [47]. In the eye-in-hand configuration, converting quantities between the two spaces without considering the viewing angle would lead to a loss of motion correspondence between the movements expected by the operator and those actually executed by the robot in the camera view. Therefore, in our EHCC, we transfer the HD’s twist to the FR through an eye-hand coordination mapping and introduce dynamic scaling factors to improve control accuracy based on the task requirements.
@@ III-B3 Eye-hand coordination
III-C Robot interaction control
III-C1 Admittance model
@@
III-C2 Free-hand force limitation strategy
@@
III-D Haptic feedback controller
In performing medical procedures, the operator relies heavily on the sense of touch. Therefore, in a HBTS providing realistic haptic sensations, it becomes critical to the success of the procedures. In impedance systems, the interaction is captured by both the contact force ${\bf f}_{e}^{\mathcal{R}_{e}}$ and the position error $\tilde{\bf p}^{\mathcal{R}_{e}}$, the latter being characterized by slower transients than the former, which exhibits a step dynamics. As already validated in a previous work [42], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks. Therefore, the proposed HFC defines a second (virtual) impedance model, generating the force feedback from the motion error of the admittance controller:
@@ IV Experimental validation
To evaluate the performance of the proposed HBTS, we conduct interaction experiments using a phantom. Each experiment involves reaching a designated target tooth with an inactive dental scaler and applying a force to its surface, keeping the tool stationary. During the experiments, the robot remains out of the human operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD. We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right first molar, as shown in Figure 6. The chosen targets are significant since each requires a different tool orientation, making the use of eye-hand coordination essential. For each target, we perform three interaction experiments with active force limitation and three without. With the aim of confirming the impact of the proposed strategy on the safety of interaction, we estimate, from the data recorded during the experiments, the conditional expectation of the interaction forces given a virtual penetration commanded by the operator. The conducted analysis is reported in Section IV-B.
Figure 6: Target teeth for dental scaling experiments: 1) lower right first premolar; 2) upper right central incisor; 3) upper right first molar.
@@ IV-A3 Admittance controller parametrization
Since the interaction experiments require modest compliance and high damping along the normal axis, and greater and more precise positioning along the tangential axes, we set the impedance parameters reported in Table I.
Axis of ${\mathcal{R}_{e}}$ ${\bf K}_{P}(N/m)$ ${\bf K}_{D}(Ns/m)$ ${\bf M}_{d}(kg)$
x $50000$ $2500$ $50$
y $50000$ $2500$ $50$
z $6000$ $12000$ $50$
TABLE I: Admittance controller parameters
@@ IV-A4 Virtual feedback parametrization
With regard to the virtualization of the force feedback, the virtual impedance matrices are tuned to provide the human operator with a natural sensation, ensuring both smooth transient response and reduced vibrations. Therefore, we set the parameters reported in Table II.
@@
Axis of ${\mathcal{H}_{b}}$ ${\bf K}_{h}(N/m)$ ${\bf D}_{h}(Ns/m)$
x $0.1$ $0.5$
y $0.1$ $0.5$
z $0.1$ $0.5$
TABLE II: Virtual feedback parameters.
@@ IV-B Force limitation assessment
Given the conditional probability function $p\left(a(k)|b(k)\right)$, we estimate the conditional expectation $E\left(A|B=b_{d}\right),\forall b_{d}\in\mathcal{B}_{d}$:
$E\left(A|B=b_{d}\right)=\sum_{a_{d}\in\mathcal{A}_{d}}a_{d}~{}p(A=a_{d}|B=b_{d}),$ (34)
where $a_{d}$ is the discrete value of a bin in $\mathcal{A}_{d}$, and the conditional variance $\text{Var}(A|B=b_{d}),\forall b_{d}\in\mathcal{B}_{d}$:
$\begin{split}\text{Var}(A|B=b_{d})=&\sum_{a_{d}\in\mathcal{A}_{d}}(a_{d}-E(A|B=b_{d}))^{2}\\ &\quad\cdot p(A=a_{d}|B=b_{d}).\end{split}$ (35)
In Figure 8 are reported the conditional expectations of the exerted force $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and the conditional standard deviation, obtained as the square root of the conditional variance, in both scenarios.
@@
To evaluate the effectiveness of the force limitation strategy, we compared $E\left(A|B=b_{d}\right),\forall b_{d}\in\mathcal{B}_{d}$ and $\text{Var}(A|B=b_{d}),\forall b_{d}\in\mathcal{B}_{d}$ in both scenarios. To test whether the conditional variances differ significantly, we perform Levene’s test and Bartlett’s test. Levene’s test shows that the conditional variances are significantly different (p-value $p\approx 0$), indicating that force limitation strategy affects the dispersion of normal force $f_{e}^{\mathcal{R}_{e},\perp}(k)$ with respect to virtual penetration $\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\right\rVert$. Bartlett’s test, which is more sensitive to normal distributions, confirms this observation ($p\approx 0$). This suggests that the difference in variance is not only due to the sampling effect but is a structural feature of the collected data. We compare the conditional expectation using a t-test for independent samples. Since Levene’s test suggests a significant variance difference, we apply Welch’s t-test, which is more robust in the case of unequal variances. The t-test returned a t-statistic value of $t=-6.18$, confirming that the normal force is lower in Scenario A, and $p\approx 0$. Since such value is below the significance threshold $\alpha=0.05$, we can reject the null hypothesis and conclude that the conditional expectations of normal force in the two scenarios are significantly different.
@@ IV-C Eye-hand coordination assessment
These results confirm that the mapping of motion between HD space and FR space is affected by the rotation of the visual field. However, they do not provide any suggestion regarding the effectiveness of eye-hand coordination for the human operator. Therefore, in we provide a comprehensive qualitative assessment of the latter.
@@ IV-D End-to-end system experimental demonstration
We provide a demonstration of the end-to-end system’s operation in the accompanying video , where we showcase the execution of the dental scaling procedure on the three target teeth in Scenario A. In this section, we compare the outcomes of the execution of a dental scaling procedure in Scenario A, against those obtained in Scenario B. To avoid repetition, the analysis focuses on the outcomes of executing a dental scaling action on a single target tooth, the lower right first premolar.
In Figure 10 the normal forces exerted by the FR in both scenarios are shown. A visual inspection showcases a significant reduction in interaction forces when the free-hand force limitation strategy is active, with a decrease of approximately $50\%$ in Scenario A, thereby confirming the validation presented in Section IV-B. The effectiveness of this strategy is made even more evident by Figure 11, which shows the norm of the references provided to the admittance controller. Notably, in Scenario A, where the interaction forces are lower, the virtual penetration commanded to the controller is even greater than in Scenario B. Nevertheless, as shown in Figure 11a, the implementation of this strategy enables the scaling of the provided references, effectively reducing the contact forces and enhancing the system’s safety.
(a) Contact forces in Scenario A.
(b) Contact forces in Scenario B.
Figure 10: Nomal forces in both scenarios.
(a) Norm of the admittance controller references in Scenario A.
(b) Norm of the admittance controller references in Scenario B.
Figure 11: Norm of the admittance controller references in both scenarios.
Finally, in Figure 12 the forces rendered on the HD in both scenarios are shown. As expected, the results are consistent between the two scenarios, while both showing high similarity to exerted force profiles as well as slow dynamics. Such aspects enhance naturalness for the human operator.
(a) Forces rendered on the HD in Scenario A.
(b) Forces rendered on the HD in Scenario B.
Figure 12: Forces rendered on the HD in both scenarios.
@@ V Discussions and conclusions
Motivated by a lack of research on robotized free-hand dental procedures, this paper proposed a HBTS specifically designed to this aim. The system features visual and haptic feedback, that are appropriately combined and coordinated to allow for a natural teleoperation experience. To improve the visibility of the intervention site, we designed a dedicated end-effector accommodating an eye-in-hand endoscopic camera, and ensuring compatibility with existing clinical tools. We improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities. By providing the eye-hand coordination and the TCP control, we enabled the human operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise. Additionally, by appropriately transforming the virtualized force feedback, we also ensured the force correspondence, enabling the human operator to experience a highly natural sensory perception. Finally, we enabled safe interactions, by scaling the motion references provided to the admittance controller. The proposed strategy allows for accurate force limitation in any contact state and for any orientation of the interaction tool, without requiring any prior modeling of the intervention site.
We believe that our HBTS is a necessary intermediate step for a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities. With respect to manual procedures, our system allows to preserve operator’s ergonomics, track operator’s performance, and enable remote healthcare.
We emphasize that the proposed system allows for the collection of haptic and motion data, which can be crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [55], or for the derivation of data-driven control policies [56] for autonomous robots.
With our future work we intend to extend the system with autonomous patient motion compensation to further enhance the system’s safety. Likewise, we intend to conduct experiments with professionals and active dental tools both to confirm the system’s usability under more representative clinical conditions and to collect expert policies for (semi-)autonomous procedures.

### v2

@@ III-A2 Human tremors filtering
III-A3 Eye-hand coordination
In a HBTS, coupling two robots with different kinematics requires the use of mapping and scaling techniques. The literature shows that this coupling can be achieved by converting the motion of the HD into that of the FR through simple kinematic transformations, and appropriately scaling it to either expand the workspace size of the HD or enhance control accuracy [46]. In the eye-in-hand configuration, converting quantities between the two spaces without considering the viewing angle would lead to a loss of motion correspondence between the movements expected by the operator and those actually executed by the robot in the camera view. Therefore, in our EHCC, we transfer the HD’s twist to the FR through an eye-hand coordination mapping and introduce dynamic scaling factors to improve control accuracy based on the task requirements.
@@ III-A3 Eye-hand coordination
III-B Robot interaction control
In this section, we first formalize the contact model between the robot and the environment (Section III-B1), and subsequently, we present the free-hand strategy for limiting the interaction forces (Section III-B2).
III-B1 Admittance model
@@
III-B2 Free-hand force limitation strategy
@@
III-C Haptic feedback controller
In performing medical procedures, the operator relies heavily on the sense of touch. Therefore, in a HBTS providing realistic haptic sensations, it becomes critical to the success of the procedures. In impedance systems, the interaction is captured by both the contact force ${\bf f}_{e}^{\mathcal{R}_{e}}$ and the position error $\tilde{\bf p}^{\mathcal{R}_{e}}$, the latter being characterized by slower transients than the former, which exhibits a step dynamics. As proposed in [49], replacing ${\bf f}_{e}^{\mathcal{R}_{e}}$ with $\tilde{\bf p}^{\mathcal{R}_{e}}$ to generate the force feedback makes each contact state transparent to the operator, preventing them from having to deal with sudden force changes they are not used to from previous training and fieldworks. Therefore, the proposed HFC defines a second (virtual) impedance model, generating the force feedback from the motion error of the admittance controller:
@@ IV Experimental validation
To evaluate the performance of the proposed FH-HBTS, we conduct interaction experiments using a phantom. Each experiment involves reaching a designated target tooth with an inactive dental scaler and applying a force to its surface, keeping the tool stationary. During the experiments, the robot remains out of the human operator’s view, who relies solely on the visual feedback from the camera and the force feedback from the HD. We select three target teeth: the lower right first premolar, the upper right central incisor, and the upper right second premolar, as shown in Figure 6. The chosen targets are significant since each requires a different tool orientation, making the use of eye-hand coordination essential. For each target, we perform three interaction experiments with active force limitation and three without. With the aim of confirming the impact of the proposed strategy on the safety of interaction, we estimate, from the data recorded during the experiments, the conditional expectation of the interaction forces given a virtual penetration commanded by the operator. The conducted analysis is reported in Section IV-B.
Figure 6: Target teeth for dental scaling experiments: 1) lower right first premolar; 2) upper right central incisor; 3) upper right second premolar.
@@ IV-A3 Admittance controller parametrization
Since the interaction experiments require modest compliance and high damping along the normal axis, and greater and more precise positioning along the tangential axes, we set ${\bf K}_{P}=\text{diag}(50000,50000,6000)$, ${\bf K}_{D}=\text{diag}(2500,2500,12000)$ and ${\bf M}_{d}=\text{diag}(50,50,50)$.
@@ IV-A4 Virtual feedback parametrization
With regard to the virtualization of the force feedback, the virtual impedance matrices are tuned to provide the human operator with a natural sensation, ensuring both smooth transient response and reduced vibrations. Therefore, we set ${\bf K}_{h}=\text{diag}(0.1,0.1,0.1)$, ${\bf D}_{h}=\text{diag}(0.5,0.5,0.5)$.
@@ IV-B Force limitation assessment
Given the conditional probability function $p\left(a(k)|b(k)\right)$, we estimate the conditional expectation $E\left(A|B\right)$, and the conditional variance $\text{Var}(A|B)$ following [55]. In Figure 8 are reported the conditional expectations of the exerted force $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and the conditional standard deviation, obtained as the square root of the conditional variance $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$, in both scenarios. Visual inspection showcases a significant reduction in the expected normal interaction force in Scenario A, given the same virtual penetration commanded.
@@
To evaluate the effectiveness of the force limitation strategy, we compare $E\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ and $\text{Var}\left(f_{e}^{\mathcal{R}_{e},\perp}\middle|\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}\right\rVert\right)$ in both scenarios. To test whether the conditional variances differ significantly, we perform Levene’s test and Bartlett’s test. Levene’s test shows that the conditional variances are significantly different (p-value $p\approx 0$), indicating that force limitation strategy affects the dispersion of normal force $f_{e}^{\mathcal{R}_{e},\perp}(k)$ with respect to virtual penetration $\left\lVert{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\right\rVert$. Bartlett’s test, which is more sensitive to normal distributions, confirms this observation ($p\approx 0$). This suggests that the difference in variance is not only due to the sampling effect but is a structural feature of the collected data. We compare the conditional expectation using a t-test for independent samples. Since Levene’s test suggests a significant variance difference, we apply Welch’s t-test, which is more robust in the case of unequal variances. The t-test returned a t-statistic value of $t=-6.18$, confirming that the normal force is lower in Scenario A, and $p\approx 0$. Since such value is below the significance threshold $\alpha=0.05$, we can reject the null hypothesis and conclude that the conditional expectations of normal force in the two scenarios are significantly different.
@@ IV-C Eye-hand coordination assessment
These results confirm that the mapping of motion between HD space and FR space is affected by the rotation of the visual field. However, they do not provide any suggestion regarding the effectiveness of eye-hand coordination for the human operator. Therefore, in the accompanying video [^1] This paper has supplementary downloadable material available at http://ieeexplore.ieee.org, provided by the authors. This includes a multimedia MOV format movie clip, showing the end-to-end system experimental demonstration. This material is 51.4 MB in size. we provide a comprehensive qualitative assessment of the latter.
@@ IV-D End-to-end system experimental demonstration
We provide a demonstration of the end-to-end system’s operation in the accompanying video, where we showcase the execution of the dental scaling procedure on the three target teeth in Scenario A. In this section, we report the outcomes of the execution of a dental scaling procedure in Scenario A, highlighting two instants of the experiment (Figure 10), and compare them against those obtained in Scenario B (Figure 11). To avoid repetition, the analysis focuses on the outcomes of executing a dental scaling action on a single target tooth, the upper right second premolar.
In Figure 10d and Figure 11b, the normal forces exerted by the FR in both scenarios are shown. A visual inspection showcases a significant reduction in interaction forces when the free-hand force limitation strategy is active, with a decrease of approximately $50\%$ in Scenario A, thereby confirming the validation presented in Section IV-B. The effectiveness of this strategy is made even more evident by comparing Figure 10c and Figure 11a. Notably, in Scenario A, where the interaction forces are lower, the virtual penetration commanded to the controller is even greater than in Scenario B. Nevertheless, as shown in Figure 10c, the implementation of this strategy enables the scaling of the provided references, effectively reducing the contact forces and enhancing the system’s safety.
[…]
(c) Norm of the admittance controller reference
(d) Contact forces
(e) Forces rendered on the HD
Figure 10: Snapshots and outcomes of the dental scaling experiment in Scenario A.
(a) Norm of the admittance controller reference
(b) Contact forces
(c) Forces rendered on the HD
Figure 11: Outcomes of the dental scaling experiment in Scenario B.
Finally, in Figure 10e and Figure 11c the forces rendered on the HD in both scenarios are shown. As expected, the results are consistent between the two scenarios, while both showing high similarity to exerted force profiles as well as slow dynamics. Such aspects enhance naturalness for the human operator.
@@ V Discussions and conclusions
This paper proposed a FH-HBTS for free-hand dental procedures. The system features visual and haptic feedback, that are appropriately combined and coordinated to allow for a natural teleoperation experience. To improve the visibility of the intervention site, we designed a dedicated end-effector accommodating an eye-in-hand endoscopic camera, and ensuring compatibility with existing clinical tools. We also improved the accuracy of the procedure by introducing human tremor filtering and dynamic scaling of velocities. By providing the eye-hand coordination and the TCP control, we enabled the human operator to perceive the motion correspondence between their movements and those of the end-effector, observed in the camera view, preserving manual dexterity and expertise. Additionally, by appropriately transforming the virtualized force feedback, we also ensured the force correspondence, enabling the human operator to experience a highly natural sensory perception. Finally, we enabled safe interactions, by scaling the motion references provided to the admittance controller. The proposed strategy allows for accurate force limitation in any contact state and for any orientation of the interaction tool, without requiring any prior modeling of the intervention site.
We believe that our FH-HBTS represents a crucial step towards a broader adoption of robotics in dentistry, as it allows for accurate treatments, while preserving the operator’s cognitive and technical abilities. With respect to manual procedures, our system allows to preserve operator’s ergonomics, track operator’s performance, and enable remote healthcare.
We emphasize that the proposed system allows to collect haptic and motion data. While this feature is crucial during the training phase of new operators for the evaluation of learning metrics, i.e learning curves [56], it is also central towards the deployment of data-driven control architectures that leverage demonstrations to synthesize control policies [57]. Finally, these data can also be used to learn cost functions [58], and hence intentions, of experts fulfilling a given task. In the context of dental procedures, this means that data collected from experts executing the procedure can be used to train policies for autonomous robots.
With our future work we plan to extend the FH-HBTS with autonomous patient motion compensation to further enhance the system’s safety, and to benchmark the performance against the da Vinci Surgical SystemTM. Also, we plan to conduct experiments with professionals and active dental tools both to confirm the system usability under more representative clinical conditions and to collect expert policies for (semi-)autonomous procedures.